A Complete Bibliography of Publications in Fundamenta Informaticae: (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

02 May 2023 Version 1.01

Title word cross-reference

(g, d, k) [25]. $-Calculus [245]. 1 [5]. b [190]. DL [83, 115]. S [82]. e [91]. k [149, 20]. K_{4} [91]. L_{1} [11]. A_{n} [198]. A_{r} [215]. K [95]. q [18]. t [211]. T_{1} [7]. × [83, 115].

-ary [211]. -Bounded [190]. -Connectivity [25]. -edge-Connectivity [25]. -exhaustions [184]. -Fault [5]. -free [91]. -Matrices [18]. -means [20]. -rendezvous [149].

1966 [228].

2-domination [213].

Aalst [184]. Abstract [95]. Abstraction [237, 103, 241]. Accepted [150, 144]. Accessibility [31]. Ackermann [183]. Action [22]. Actions [107]. Active [178]. Activity [11]. Actors [94]. Acyclic [205]. Adaptive [247, 30]. Adding [157]. Addition [243]. Additional [66]. Address [244]. Affine [192, 223]. Again [212]. agent [59]. Aggregated [203]. Aggregates [80]. Aggregation [246]. Algebra [132, 1, 129]. Algebraic [51, 151, 86, 129]. Algebras [223, 85]. Algorithm [199, 89, 220, 20]. Algorithmic [239]. Algorithms [84, 177, 88, 245, 35, 110, 46, 152, 9, 179]. Allowing [239]. Alternative [200]. Analyses [172]. Analysis [80, 171, 127, 88, 81, 131]. Analyzing [53]. and/or [97]. angle [12]. Answer [96, 78]. Antimagic [211, 45]. Anywhere [147]. Apartness [204]. Application [192, 237]. Applications [217, 186, 46, 200]. Approach [117, 83, 9, 146, 78, 176, 12, 136]. Approximate [177]. Approximations [125, 130]. Arc [55]. Arrays [40]. Articulations [186]. ary [211]. Assessments [156]. Association [114]. Associativity [97]. Attack [172]. Attack-fault [172]. Attenuation [11]. Augmented [71]. Authenticated [27]. Autoepistemic [78]. Automata [248, 150, 209, 138, 63, 197, 159, 110, 152, 141, 161, 233, 106, 144]. Automated [125, 195, 15]. Automath [218]. Auxiliary [141]. Aware [170, 26, 43]. aWide [89]. Axiomatizing [72]. Axioms [97].

Back [212]. Baire [225]. Balanced [109]. Balancing [118]. Band [118]. Barnette [242]. Base [48]. Based [237, 175, 197, 195, 47, 37, 115, 53]. Basic [239]. BCK [47]. BCK-codes [47]. Behavioural [84]. between [164]. Betweenness [177]. BGM [220]. BGN [246]. Bijective [199]. Binary [223, 46, 8, 23]. Binding [30]. Bipartite [38]. Bisimilarity [31]. Boolean [127, 156, 52, 161, 182]. Bound [149]. Bounded [190]. Bounds [86]. BPP [145]. Brain [10]. Branching [207]. Breaker [210]. Bruijn [218]. Buchberger [199]. Büchi [152]. Buffer [105]. Built [244]. Built-in [244]. Butterfly [180]. Byzantine [27].

Calculi [229]. Calculus [90]. Call [192]. Cancer [81]. Capacity [143]. Cartan [167]. Cartesian [109]. Case [156]. Causal [33, 145, 111]. Causal-Consistent [111]. Cause [122]. Cause-Effect [122]. Cellular [209, 233]. Centers [192]. Centrality [177, 10]. centric [57]. Change [247]. Characterisation [105]. Characterisations [151]. Characterising [10]. Characteristics [218]. Check [47]. Checker [116, 218]. Checking [237, 102, 114, 36, 196]. choice [184, 165]. CIFF [82]. Circuit [53]. Circular [163]. Classification [85, 16, 198, 167]. Classroom [44]. Clinical [74, 54, 176]. Clock [106]. Clock-Dependent [106]. Closed [100]. Closeness [66]. Closure [79, 119]. Closures [248]. Cluster [70]. Clustering [203]. Clusters [165]. codes [47]. Coinductive [152]. Coloring [211, 45]. Column [26]. Combinations [97]. Combinatorial [46]. Combining [36, 41]. Communication [69]. Commutativity [97]. Commuting [18]. Compactness [4]. Companion [167]. Companions [64]. Comparative [125]. Comparing [58]. Comparison [51, 75]. Comparisons [21]. Compilation [53]. Complementary [112]. Complete [109, 211, 230, 134, 141]. Completeness [223, 245]. Complex [123, 30]. Complexity [102, 156, 239, 238, 162, 35, 142, 190, 160]. Composition [44, 161]. Compositional [30]. Compression [166, 200]. Computable [183]. Computation [87, 54]. Computational [85, 219, 160]. Computationally [134]. Computations [205]. Computers [163]. Computing [38, 177, 206, 240, 128, 68, 92, 173, 176]. Concept [131]. Concurrent [236, 98, 111]. Conditional [162, 25]. Conditional- [25]. Conditions [170]. Congress [228]. Congruence [197]. Congruence-Based [197]. Connected [5]. Connectivity [25, 10]. Consensus [27]. Conservation [233]. Consistency [20, 55]. Consistent [21, 111]. Constraint [98, 35, 99, 55, 181]. Constraints [171, 195]. Constructing [71]. Construction [207]. Constructive [204]. Constructor [37]. Constructor-Based [37]. Consumption [94]. Consumption-Preserving [94]. Context [134]. Contextual [143]. Contingent [43]. Continuous [131]. Contract [95, 36]. Control [164, 171]. Cooperative [94]. Copyful [104]. Core [80]. Corrector [19, 89]. Correlations [22]. Cost [248, 245, 189, 22]. Counters [159]. Counting [153]. Coverability [187]. Coxeter [215]. CP [184]. Critical [52, 234]. Crossover [51]. Crystal [202]. CSP [40]. Cubes [71]. Cubicle [116]. Curry [36]. Curve [147]. Curves [119, 200]. Cycle [109]. Cycles [213, 242]. Cyclic [205]. Cylinders [213].

D4 [83, 115]. Daniel [231]. Data [74, 170, 246, 75, 173, 81]. Data-Aware [170]. Databases [26]. DBpedia [16]. Decentralized [168]. Decidability [243, 129, 140]. Decidable [156]. Decision [120]. Declarative [116]. Deep [174]. Definability [243]. Definite [167]. Degenerate [58]. degree [200]. Dependent [106]. Deriving [100]. Description [83, 115, 136]. Descriptional [162, 219, 142]. Descriptions [226]. Detailed [175]. Detection [168]. Determine [75]. Deterministic [2, 201, 194, 229]. Detour [5]. Development [204]. Diagnosability [91]. Diagnosis [91, 220, 168]. Diagnosis/prediction [168]. Diamond [108]. Different [75]. Digital [13, 54, 119]. Dimension [87]. dimensional [233]. Directed [177, 24]. Dischargeable [82]. Discovering [57]. Discovery [17]. Discrete [92, 168]. Discrete-event [168]. Disease [75]. Disjunctive [39]. Dissimilar [75]. Distributed [123, 116, 175]. Domains [131]. Domination [39, 214, 41, 213]. Downward [248]. Dynamic [59, 98, 35, 36]. Dynamical [192, 146]. Dynkin [198, 215].

early [218]. Edge [38, 180, 25]. EEG [74]. Effect [122]. Efficiency [123]. Efficient [87, 17, 9, 90, 179, 100]. Element [206]. Elementary [122]. Elements [128]. Elliptic [200]. Embedding [97, 142]. Emergency [192]. Emerging [222]. Emission [11]. Encoding [127]. Energy [202]. Ensemble [126, 70]. Entries [46]. environments [238]. Epileptic [74]. Epistemic [224]. Equal [46]. Equations [88]. Equivalence [31]. Essay [15]. Essential [64]. Estimation [7, 11]. Euler [2]. Evaluations [45]. Event [123, 168]. Events [193]. Evolving [114]. Exact [177]. Exceptions [79]. exhaustions [184]. Experimental [51]. Exploration [59]. Exploring [9]. Expressions [216, 124]. Expressive [194]. Extensions [122, 199]. Extra [72].

Face [45]. factor [234]. Factorization [2]. Factors [153]. Fair [107]. Faithful [69]. Falsity [227]. Fault [52, 5, 168, 172]. Fault-Tolerant [52]. Feature [175]. Features [157]. Field [206, 12]. Finding [242, 40]. Finite [49, 150, 63, 197, 110, 219, 48]. Finite-State [219]. Finitely [227]. Finitely-many [227]. Finiteness [187]. First [160]. First-order [160]. Fixed [49]. Fixpoint [102]. Flash [26]. Flash-Aware [26]. Flexible [61]. Flight [171]. Flow [233]. Forcing [180]. Formal [94, 131]. Forms [198, 215, 100]. Foundations [204]. Fractional [234]. Fragments [156]. Framework [82, 95, 7, 198]. Free [223, 134, 184, 165, 178, 91]. Free-choice [184, 165]. Freeness [135]. Frozen [107]. Full [211]. Function [2, 183]. Functional [10]. Functions [24, 68, 200, 182].

Games [225, 210]. Gathering [232]. Gaussian [76]. Gene [127]. General [112]. Generalization [227, 234]. Generalized [88, 125, 120]. Generation [217, 146]. Generative [143]. Generator [244]. Generators [100]. Geometrically [203]. Geometrically-Modeled [203]. Getting [212]. Global [206]. Graders [15]. Grading [15]. Gram [198]. Grammars [164, 162, 134, 3, 142, 143]. Granular [126, 176]. Granulation [173]. Graph [108, 87, 198]. Graph-theoretic [87]. Graphs [38, 109, 45, 133, 177, 39, 242, 24, 230, 91, 214, 179, 5, 234]. green [163]. Grey [13]. Grid [232]. Grids [72]. Grouping [15]. Groups [75]. Growth [66].

Hamiltonian [242]. Handling [35]. Hardness [202]. Hardware [30]. Hardware-software [30]. Hashing [175]. Haskell [94]. Hennessy [33]. Heuristic [53]. Heuristic-based [53]. High [200]. High-degree [200]. Higher [102, 24, 216]. Higher-Order [102, 24]. Hitting [207]. Hoare [132]. Home [165]. Homeomorphic [97]. Hospital [176]. Human [43]. Human-aware [43]. Hybrid [27]. Hypercube [220]. Hypercube-like [220]. Hypercubes [25].

ID [234]. ID-factor-critical [234]. Identifying [74]. Images [7, 13]. Immediate [249]. Implementations [35]. Improved [115, 162]. Improvement [15]. Improving [16]. Impure [238]. Incomplete [120]. Independent [71, 178]. Index [10, 130]. Indiscernibility [130]. Induced [179]. Inductive [87]. Infeasible [89, 19]. Inference [110]. Inferring [193]. Infinite [232, 225, 194, 130]. Information [16, 173, 176]. Insecure [246]. Insertion [140, 3]. Integer [2]. Interference [155]. Interior [89, 19]. Interior-point [89, 19]. International [228]. Interpolation [9].

Intersection [166]. Interval [163, 92]. Interval-valued [163]. Introduction [236]. Invariance [8]. Invariants [240, 215]. Inverse [183]. Investigating [188]. Irregular [133]. Irregularity [38]. issue [42]. Issues [243, 239, 238]. Iterated [219].

Jacobi [166]. Joint [7, 11]. Jordan [119]. Jump [92]. Justifications [35].

Kernel [217]. KNN [174]. Knowledge [157].

Labeling [61]. Labelled [3]. Lambda [100]. Landscape [222]. Language [153, 201]. Languages [157, 98, 63, 151, 139, 3, 130, 143, 144]. Large [87, 17]. Launcher [171]. Learning [90]. Length [206, 150]. Let [216]. Letter [65]. Letter-to-Letter [65]. level [117]. Lexicographic [79, 214]. like [220]. Likelihood [7]. Limited [12]. Limited-angle [12]. Linear [149, 46, 89, 118, 183]. Link [17, 41, 181]. Liveness [249]. Load [118]. Local [211, 45, 220, 195]. Locality [175]. Locally [139]. Locating [40]. Logarithm [92]. Logic [224, 132, 102, 33, 83, 115, 49, 32, 99, 37, 78, 81, 160]. Logical [35, 151]. Logics [31, 227, 229]. Loss [174]. Lower [86]. Lowest [22]. lpopt [96]. LR [46]. LR-tableaux [46]. Lucent [184, 165].

Machine [90]. Machines [163]. Magnetic [7]. Many [31, 229, 32, 227]. Many-sorted [32]. Many-valued [31, 229]. Mapping [61]. Matching [230, 179, 216]. Mathematicians [228]. Matrices [18, 149, 21, 167, 135]. Matrix [47, 48, 135]. Matroids [136]. Maximum [7, 179]. Means [87, 20]. Mechanisms [164]. Meeting [232]. Membranes [178]. memoir [228]. Memory [247, 244]. Mereological [4]. Merging [247]. Message [111]. Meta [245]. Meta-Search [245]. Metalogic [157]. Method [19, 48, 92, 168]. Methods [75, 41]. Metric [49]. Metrics [245]. Micro [228]. Micro-memoir [228]. Milner [33]. MILP [52]. Minimal [159, 110, 48]. Minimisation [202]. Minimization [11]. Minimum [40, 181]. Mining [74, 75, 176]. Mission [52]. Modal [31, 32]. Model [80, 237, 126, 102, 61, 116, 203, 91, 220, 30, 196]. Model-Checking [102, 196]. Modeled [203]. Models [192, 117, 170, 86, 195, 140, 200]. Modulo [97, 128]. Monadic [226]. Mood [76]. Morphisms [208, 23]. Motion [7, 64]. Moves [145]. Multi [59, 117, 105]. Multi-agent [59]. Multi-Buffer [105]. Multi-level [117]. Multiagent [42]. Multicore [118]. Multicounter [159]. Multigranular [67]. Multilinear [86]. Multipartite [109]. Multiset [201]. Music [76].

negative [198, 215]. Neighborhood [89, 19]. Neighbourhood [55]. Net [60, 186, 239]. Nets [192, 237, 205, 62, 238, 188, 240, 187, 194, 145, 249, 189, 190, 184, 57, 165]. Network [1, 10, 70]. Networks [127, 133, 52, 66, 220, 161, 180, 41]. Nilpotent [46]. no [72]. Node [71]. Node-Independent [71]. Nodes [232, 52]. Nominal [216]. Non [194, 198, 215, 155, 195, 229, 72, 134, 142, 201]. Non-Context-Free [134]. Non-Deterministic [201, 194, 229]. Non-Interference [155]. Non-local [195]. Non-negative [198, 215]. Non-Self-Embedding [142]. Non-unary [72]. Nondeterministic [110, 219]. Nonmonotonic [78]. Nonregular [63]. norm [19]. Normal [100]. Note [229, 234]. Novel [126]. NTRU [88]. Number [217, 39, 213, 112, 5, 233]. Numerical [68]. NZ [149].

Obituary [231]. Object [57]. Object-centric [57]. Obligations [82]. Observation [249]. OMT [52]. Once [207]. One [108, 19, 134, 233]. One-dimensional [233]. One-norm [19]. One-Rule [108]. onto [109]. Ontologies [79]. Ontology [16]. Operations [46]. Operator [119]. Operators [51, 46]. Optimal [109, 117, 52, 54]. Optimization [96, 89, 19, 70]. Optimum [80]. Oracles [2]. Order [97, 87, 102, 226, 24, 216, 19, 160]. Order- [87]. Order-sorted [97]. Order-theoretic [226]. Organisms [30]. Oriented [26, 60]. Orthogonal [21].

P [178, 68]. Pairwise [21]. Palindromic [208]. Parallel [110, 100]. Parameter [103, 9]. Parameterized [116, 240, 86]. Parameters [7]. Parametric [172, 171, 189, 29]. Parikh [18, 135]. Parity [47]. Parkinson [75]. Partial [55]. Particle [233]. Passing [111]. Path [109]. Paths [181]. Patient [54]. Patients [74, 75]. Pattern [74]. PBW [199]. Perfect [214]. Periodic [59]. Permutation [51, 179]. Perpetual [184]. Persistent [155]. Personalised [54]. Perspective [197]. Petri [192, 237, 205, 60, 62, 186, 239, 238, 188, 240, 187, 194, 249, 189, 190, 184, 57]. Phase [247, 12]. Physical [160]. Piecewise [192]. Place [238, 147]. Place-environments [238]. Placements [52]. Plane [45]. Planning [43]. PMC [91]. point [89, 19]. Points [203, 181]. Polyadic [32]. Polygons [181]. Polyhedral [237]. Polynomial [167, 207]. Polynomial-Time [207, 167]. Polynomials [230]. Polyominoes [146]. Position [112]. Positive [167]. Positron [11]. Power [222, 194, 219]. Powers [63]. Practice [204]. Pre [129]. Pre-rough [129]. Precision [120]. Prediction [202, 168]. Predictor [89]. Predictor-corrector [89]. Preface [42, 221, 14, 6, 235, 154, 185, 158, 50, 113, 34, 73, 101, 93, 137, 191, 169, 28, 77, 121, 148, 173, 56]. Preserving [94, 208, 246]. Primitive [149]. Prisms [112]. Privacy [246]. Probabilistic [236, 117, 106]. Probabilities [106]. Problem [178, 44, 179, 53, 92, 135]. Problems [51, 189, 55]. Process [132, 61, 76, 170, 195, 176]. Processing [123]. Processors [118]. Product [109, 214]. Production [134]. Products [186]. Programming [96, 99, 78, 81]. Programs [103, 111, 207]. Progression [75]. Projection [8]. Projections [21]. Proof [184, 218]. Properties [63, 241, 182]. Property [4]. Propositional [49]. Protocol [27]. Protocols [116, 246]. Proving [86]. Pseudo [52]. Pseudo-Boolean [52]. Pure [114, 238]. Pushdown [138]. Pythagoras [206].

Quantitative [124]. Quantum [53]. Quasi [167]. Quasi-Cartan [167]. Query [181, 201]. Query-Points [181]. Questions [140]. Queue [138]. Queues [193].

Random [217, 126]. Randomized [53]. Ranges [49]. Rank [24]. Reachability [37]. Reaction [127, 64]. Reactivity [171]. Read [207]. Read-Once [207]. Real [243, 103, 159]. Real-Time [103, 159]. Reasoner [115]. Reasoning [123, 33, 83, 79]. Recognizing [133, 63]. Reconstruction [12]. Rectangular [72]. Recursive [84, 102, 187, 216]. Red [163]. Reduction [108]. Refinement [117, 103]. Reflections [126]. Reflexive [78]. Region [195]. Regression [76]. Regular [63, 151, 124]. Regularity [226]. Regularized [9]. Regulatory [127]. Related [38, 140]. Relations [164, 31, 72, 69, 130]. Relaxation [61]. Relay [52]. rendezvous [149]. Repair [195]. Replay [111]. Representation [157]. Residual [41]. Residuated [129]. resolution [7]. Resonance [7]. Resource [94]. Resources [193]. Respect [119]. Restarting [141, 144]. Restricted [238, 92]. Results [13, 162]. Retraction [35]. Reversibility [188]. Reversible [151, 111]. Reversing [205]. Rewrite [37]. Rewriting [108]. Richness [208]. Right [199]. Robust [11]. Roman [214]. Roots [63]. Rough [175, 125, 67, 120, 130, 136, 23, 129]. Rule [108, 96, 74]. Rules [114, 35, 22].

S4F [78]. Safe [248]. Safety [33]. Satisfaction [55]. Scalable [173]. Scale [13, 17, 8]. Scattered [134]. Schedulability [171]. Scheduling [94, 118]. Schemes [248]. Science [173]. Search [80, 245, 53]. Second [226, 19]. Second-order [226, 19]. Selection [175, 8, 143, 70]. Selective [99]. Self [209, 138, 244, 142]. Self-stabilisation [209]. Self-test [244]. Self-Verifying [138]. Semantics [145, 111]. Semi [162]. Semi-Conditional [162]. Semigroups [204, 135]. Sensitive [175]. Separating [153]. Sequence [244]. Sequential [164, 100]. Services [83]. Set [96, 83, 156, 115, 175, 178, 207, 78, 22, 136, 23]. Set-based [115]. Set-theoretic [83]. Sets [149, 114, 67, 120, 107]. Setting [132]. Shared [193]. Short [16]. Shortest [150]. Shuffling [65]. Sided [139]. Signatures [74]. Silent [145]. Silico [54]. Similar [15]. Similarities [17]. Simple [162, 90, 181]. Simplify [147]. Simply [100]. Simply-Typed [100]. Simson [231]. Simulation [7, 105, 163]. Singleton [55]. Singular [128]. Size [159, 86]. Skeleton [241]. Skeptical [79]. Skew [199, 167]. Skew-symmetrizable [167]. Slicing [98]. Small [213, 68]. SMT [237]. SMT-Based [237]. Social [10, 70]. software [30]. Solution [178]. Solve [27]. Solver [40]. Solvers [52]. Solving [225, 44, 118]. Some [223, 239, 125]. sorted [97, 32]. Soundness [170]. Space [225, 17, 21, 9, 48]. Spaces [4]. Spanning [71, 210]. Special [42]. Squares [206, 128]. stabilisation [209]. Stable [80]. State [195, 219, 168]. State-Based [195]. Static [36, 118]. Steps [188]. Stochastic [155, 124]. Stop [62]. Stop-transitions [62].

Storage [88, 26]. Strategies [53]. Strategy [11]. Streaming [104]. Strength [38]. Strictly [139]. String [108, 104, 17]. Strings [58, 150]. Strong [198]. Structural [117, 249, 196]. Structure [222, 202, 67]. Structured [76]. Structures [122, 210, 129]. Stubborn [107]. Study [7, 175, 125]. Subdivision [39]. Subgraphs [108]. Subregular [143]. Substitution [65]. Sum [206]. Super [7]. Super-resolution [7]. Surfaces [119]. Symbolic [95, 196]. Symbols [141]. Symmetric [19]. symmetrizable [167]. Syntactic [224]. Synthesis [95, 117, 60, 103, 186, 239, 238, 29, 190]. System [108, 171, 64, 146, 176]. Systems [236, 42, 59, 127, 186, 193, 178, 68, 140, 118, 29, 168]. Szilard [3].

tableaux [46]. Tables [120]. Tail [102]. Tail-Recursive [102]. Target [60]. Target-oriented [60]. Tasks [59]. Taxonomy [156, 141]. Team [44]. TEAMAS [42]. Teams [42]. Technique [175]. Techniques [1, 239]. Temporal [241]. Tensor [90]. Termination [187]. Terms [100]. test [244]. Testable [139]. Testing [107, 182]. Text [16]. their [88, 186, 200]. Theorem [184]. theoretic [87, 83, 226]. Theoretical [198]. Theories [37]. Theory [243, 84, 156, 35, 195, 204, 23]. There [212]. Threshold [127]. Thresholds [68]. Tilings [209]. Time [103, 149, 159, 178, 17, 189, 207, 183, 167]. Time-free [178]. Timed [59, 106]. Tolerant [52]. Tomographic [12]. Tomography [9, 8, 11]. Tool [96, 81]. Topological [105, 48]. Topology [69]. Total [39]. Totient [2]. Trace [103]. Transducers [104, 219]. Transition [186]. transitions [62]. Translation [94]. Transtopic [10]. Transversal [136]. Trapezoid [179]. Treatment [54]. Tree [197]. Trees [172, 109, 211, 71, 226]. Trials [54]. Triangular [118]. Triangulated [133]. Tubular [85]. Turing [163]. Tutte [230]. Twins [54]. Twisted [166]. Two [150, 139]. Two-Sided [139]. Two-way [150]. Type [198, 215]. Typed [100].

Unambiguous [194]. unary [72]. Uncertainty [203]. Unification [99, 216]. Unified [168]. Uniform [219]. Unifying [78]. Uniqueness [13]. Unit [198, 215]. Universal [68, 244, 241]. Unobserved [193]. Unsatisfiable [80]. Uses [246]. Using [87, 132, 195, 40, 8, 184, 61, 103, 16, 178, 17].

Valued [227, 31, 229, 163]. Variable [170, 120]. Variable-to-Variable [170]. Verification [95, 1, 103, 116, 170, 29, 124]. Verifying [138]. Vertex [211]. Vertices [147]. via [116, 233]. Vicinity [129]. Virtual [30]. Visibility [133, 181].

Walker [210]. way [150]. Weighted [144]. Wheel [38]. wide [19]. Width [207]. Wirelength [109]. Wireless [52]. without [141]. Word [65]. Words [194, 153].

References

Brodo:2020:VTN

[1] Linda Brodo and Carlos Olarte. Verification techniques for a network algebra. Fundamenta Informaticae, 172(1):1–38, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hittmeir:2020:DIF

[2] Markus Hittmeir and Jacek Pomykała. Deterministic integer factorization with oracles for Euler’s totient function. Fundamenta Informaticae, 172 (1):39–51, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paul:2020:SLL

[3] Prithwineel Paul. On Szilard languages of labelled insertion grammars. Fundamenta Informaticae, 172(1):53–72, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2020:CPM

[4] Lech Polkowski. On the compactness property of mereological spaces. Fundamenta Informaticae, 172(1):73–95, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Raghu:2020:DNF

[5] T. Venkata Raghu, R. Sundara Rajan, A. Ramesh Babu, and S. Anil. Detour number of 1-fault connected graphs. Fundamenta Informaticae, 172(1):97–104, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brunetti:2020:P

[6] Sara Brunetti, Paolo Dulio, Andrea Frosini, and Grzegorz Rozenberg. Preface. Fundamenta Informaticae, 172(2):i–xi, ???? 2020. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Beirinckx:2020:JML

[7] Quinten Beirinckx, Gabriel Ramos-Llordén, Ben Jeurissen, Dirk H. J. Poot, Paul M. Parizel, Marleen Verhoye, Jan Sijbers, and Arnold J. den Dekker. Joint maximum likelihood estimation of motion and T_{1} parameters from magnetic resonance images in a super-resolution framework: a simulation study. Fundamenta Informaticae, 172(2):105–128, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leko:2020:SIP

[8] Gábor Lékó and Péter Balázs. Scale invariance in projection selection using binary tomography. Fundamenta Informaticae, 172(2):129–142, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lagerwerf:2020:EIA

[9] Marinus J. Lagerwerf, Willem Jan Palenstijn, Folkert Bleichrodt, and K. Joost Batenburg. An efficient interpolation approach for exploring the parameter space of regularized tomography algorithms. Fundamenta Informaticae, 172(2):143–167, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Menegaz:2020:CFB

[10] Gloria Menegaz, Claudio Tomazzoli, Matteo Cristani, Ilaria Boscolo Galazzo, and Silvia Francesca Storti. Characterising functional brain connectivity as social network: the transtopic centrality index. Fundamenta Informaticae, 172(2):169–186, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Presotto:2020:MSR

[11] Luca Presotto. A L_{1} minimization strategy for robust joint activity and attenuation estimation in positron emission tomography. Fundamenta Informaticae, 172(2):187–202, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Turpin:2020:PFA

[12] Léonard Turpin, Stéphane Roux, Olivier Caty, and Sébastien Denneulin. A phase field approach to limited-angle tomographic reconstruction. Fundamenta Informaticae, 172(2):203–219, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brunetti:2020:URG

[13] Sara Brunetti, Paolo Dulio, and Carla Peri. Uniqueness results for grey scale digital images. Fundamenta Informaticae, 172(2):221–238, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Badica:2020:P

[14] Costin Badica, Mirjana Ivanovic, Yannis Manolopoulos, Riccardo Rosati, and Paolo Torroni. Preface. Fundamenta Informaticae, 172(3):i–ii, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zupanc:2020:IAE

[15] Kaja Zupanc and Zoran Bosnić. Improvement of automated essay grading by grouping similar graders. Fundamenta Informaticae, 172(3):239–259, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Flisar:2020:IST

[16] Jernej Flisar and Vili Podgorelec. Improving short text classification using information from DBpedia ontology. Fundamenta Informaticae, 172(3): 261–297, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Karampelas:2020:TSE

[17] Andreas Karampelas and George A. Vouros. Time and space efficient large scale link discovery using string similarities. Fundamenta Informaticae, 172(3):299–325, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bera:2020:CPM

[18] Somnath Bera and Kalpana Mahalingam. On commuting Parikh qmatrices. Fundamenta Informaticae, 172(4):327–341, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kheirfam:2020:SOC

[19] Behrouz Kheirfam. A second-order corrector infeasible interior-point method with one-norm wide neighborhood for symmetric optimization. Fundamenta Informaticae, 172(4):343–359, ???? 2020. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Klopotek:2020:CMA

[20] Mieczysław A. Kłopotek. On the consistency of k-means++ algorithm. Fundamenta Informaticae, 172(4):361–377, ???? 2020. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koczkodaj:2020:OPS

[21] Waldemar W. Koczkodaj, Ryszard Smarzewski, and Jacek Szybowski. On orthogonal projections on the space of consistent pairwise comparisons matrices. Fundamenta Informaticae, 172(4):379–397, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tzacheva:2020:ARL

[22] Angelina A. Tzacheva, Ramya A. Shankar, Sridharan Ramachandran, and Arunkumar Bagavathi. Action rules of lowest cost and action set correlations. Fundamenta Informaticae, 172(4):399–412, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zheng:2020:MBR

[23] Tingting Zheng. Morphisms in binary rough set theory. Fundamenta Informaticae, 172(4):413–434, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kashiwabara:2020:HOR

[24] Kenji Kashiwabara, Ikumi Horie, and Kazunori Yamaguchi. Higher-order rank functions on directed graphs. Fundamenta Informaticae, 173(1):1– 31, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:2020:CCC

[25] Cheng-Kuan Lin, Liang Ma, Jianxi Fan, Lih-Hsing Hsu, and Yuan-Hsiang Teng. The conditional-(g, d, k)-connectivity and conditional-(g, d, k)-edge-connectivity on the hypercubes. Fundamenta Informaticae, 173(1):33– 45, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Macyna:2020:FAS

[26] Wojciech Macyna and Michal Kukowski. Flash-aware storage of the column oriented databases. Fundamenta Informaticae, 173(1):47–72, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tioura:2020:HPS

[27] Abdelhamid Tioura, Hamouma Moumen, Hamoudi Kalla, and Ahmed Ait Saidi. A hybrid protocol to solve authenticated Byzantine consensus. Fundamenta Informaticae, 173(1):73–89, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kleijn:2020:P

[28] Jetty Kleijn, Laurenţiu Leuştean, and Dorel Lucanu. Preface. Fundamenta Informaticae, 173(2–3):i–ii, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sofronie-Stokkermans:2020:PSV

[29] Viorica Sofronie-Stokkermans. Parametric systems: Verification and synthesis. Fundamenta Informaticae, 173(2–3):91–138, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paduraru:2020:AVO

[30] Ciprian Ionut Paduraru and Gheorghe Stefanescu. Adaptive virtual organisms: A compositional model for complex hardware-software binding. Fundamenta Informaticae, 173(2–3):139–176, ???? 2020. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Diaconescu:2020:MEB

[31] Denisa Diaconescu. Modal equivalence and bisimilarity in many-valued modal logics with many-valued accessibility relations. Fundamenta Informaticae, 173(2–3):177–189, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leustean:2020:MSP

[32] Ioana Leuştean, Natalia Moangă, and Traian Florin Şerbănuţă. A manysorted polyadic modal logic. Fundamenta Informaticae, 173(2–3):191– 215, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Caltais:2020:CRS

[33] Georgiana Caltais, Mohammad Reza Mousavi, and Hargurbir Singh. Causal reasoning for safety in Hennessy Milner logic. Fundamenta Informaticae, 173(2–3):217–251, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fioravanti:2020:P

[34] Fabio Fioravanti, John P. Gallagher, and Maurizio Proietti. Preface. Fundamenta Informaticae, 173(4):i–ii, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fruhwirth:2020:JCH

[35] Thom Frühwirth. Justifications in constraint handling rules for logical retraction in dynamic algorithms: Theory, implementations, and complexity. Fundamenta Informaticae, 173(4):253–283, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hanus:2020:CSD

[36] Michael Hanus. Combining static and dynamic contract checking for Curry. Fundamenta Informaticae, 173(4):285–314, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skeirik:2020:CBR

[37] Stephen Skeirik, Andrei Stefanescu, and José Meseguer. A constructorbased reachability logic for rewrite theories. Fundamenta Informaticae, 173(4):315–382, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ahmad:2020:CEI

[38] Ali Ahmad, Muhammad Ahsan Asim, Basem Assiri, and Andrea Semaničová-Feňovčíková. Computing the edge irregularity strength of bipartite graphs and wheel related graphs. Fundamenta Informaticae, 174 (1):1–13, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciftci:2020:DTD

[39] Canan Çiftçi and Vecdi Aytaç. Disjunctive total domination subdivision number of graphs. Fundamenta Informaticae, 174(1):15–26, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Konishi:2020:FML

[40] Tatsuya Konishi, Hideharu Kojima, Hiroyuki Nakagawa, and Tatsuhiro Tsuchiya. Finding minimum locating arrays using a CSP solver. Fundamenta Informaticae, 174(1):27–42, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Turaci:2020:CML

[41] Tufan Turacı. On combining the methods of link residual and domination in networks. Fundamenta Informaticae, 174(1):43–59, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Andrejczuk:2020:SIT

[42] Ewa Andrejczuk, Juan M. Alberola, Leandro Marcolino, and Paolo Torroni. Special issue of teams in multiagent systems (TEAMAS): Preface. Fundamenta Informaticae, 174(1):61–62, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Andres:2020:HAC

[43] Ignasi Andrés, Leliane Nunes de Barros, and Karina Valdivia Delgado. Human-aware contingent planning. Fundamenta Informaticae, 174(1):63– 81, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Manya:2020:STC

[44] Felip Manyà, Santiago Negrete, Carme Roig, and Joan Ramon Soler. Solving the team composition problem in a classroom. Fundamenta Informaticae, 174(1):83–101, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bong:2020:LFA

[45] Novi Bong, Martin Bača, Andrea Semaničová-Feňovčíková, Kiki A. Sugeng, and Tao-Ming Wang. Local face antimagic evaluations and coloring of plane graphs. Fundamenta Informaticae, 174(2):103–119, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kaniecki:2020:CAB

[46] Mariusz Kaniecki and Justyna Kosakowska. Combinatorial algorithms for binary operations on LR-tableaux with entries equal to 1 with applications to nilpotent linear operators. Fundamenta Informaticae, 174(2):121–136, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Keshavarzian:2020:BCB

[47] Nazanin Keshavarzian, Arsham Borumand Saeid, and Abolfazl Tehranian. BCK-codes based on a parity check matrix. Fundamenta Informaticae, 174(2):137–165, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:2020:MBF

[48] Yidong Lin, Jinjin Li, Liangxue Peng, and Ziqin Feng. Minimal base for finite topological space by matrix method. Fundamenta Informaticae, 174 (2):167–183, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Djordjevic:2020:PML

[49] Radosav Djordjević, Nebojša Ikodinović, and Nenad Stojanović. A propositional metric logic with fixed finite ranges. Fundamenta Informaticae, 174(2):185–199, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eiter:2020:P

[50] Thomas Eiter, Marco Maratea, and Mauro Vallati. Preface. Fundamenta Informaticae, 174(3–4):i–iii, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baioletti:2020:ECA

[51] Marco Baioletti, Gabriele Di Bari, Alfredo Milani, and Valentino Santucci. An experimental comparison of algebraic crossover operators for permutation problems. Fundamenta Informaticae, 174(3–4):201–228, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chen:2020:MPB

[52] Qian Matteo Chen, Alberto Finzi, Toni Mancini, Igor Melatti, and Enrico Tronci. MILP, pseudo-Boolean, and OMT solvers for optimal faulttolerant placements of relay nodes in mission critical wireless networks. Fundamenta Informaticae, 174(3–4):229–258, ???? 2020. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Oddi:2020:AHB

[53] Angelo Oddi and Riccardo Rasconi. Analyzing heuristic-based randomized search strategies for the quantum circuit compilation problem. Fundamenta Informaticae, 174(3–4):259–281, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sinisi:2020:OPT

[54] Stefano Sinisi, Vadim Alimguzhin, Toni Mancini, Enrico Tronci, Federico Mari, and Brigitte Leeners. Optimal personalised treatment computation through in silico clinical trials on patient digital twins. Fundamenta Informaticae, 174(3–4):283–310, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wallace:2020:PNS

[55] Richard J. Wallace. Partial (neighbourhood) singleton arc consistency for constraint satisfaction problems. Fundamenta Informaticae, 174(3–4): 311–344, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

terBeek:2020:P

[56] Maurice ter Beek, Maciej Koutny, and Grzegorz Rozenberg. Preface. Fundamenta Informaticae, 175(1–4):v–viii, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanderAalst:2020:DOC

[57] Wil M. P. van der Aalst and Alessandro Berti. Discovering object-centric Petri nets. Fundamenta Informaticae, 175(1–4):1–40, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alzamel:2020:CDS

[58] Mai Alzamel, Lorraine A. K. Ayad, Giulia Bernardini, Roberto Grossi, Costas S. Iliopoulos, Nadia Pisanti, Solon P. Pissis, and Giovanna Rosone. Comparing degenerate strings. Fundamenta Informaticae, 175(1–4):41– 58, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Arcile:2020:DEM

[59] Johan Arcile, Raymond Devillers, and Hanna Klaudel. Dynamic exploration of multi-agent systems with periodic timed tasks. Fundamenta Informaticae, 175(1–4):59–95, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Best:2020:TOP

[60] Eike Best, Raymond Devillers, Evgeny Erofeev, and Harro Wimmel. Target-oriented Petri net synthesis. Fundamenta Informaticae, 175(1–4): 97–122, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Carmona:2020:FPM

[61] Josep Carmona, Lluís Padró, and Luis Delicado. Flexible process model mapping using relaxation labeling. Fundamenta Informaticae, 175(1–4): 123–141, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Desel:2020:STP

[62] Jörg Desel and Marc Finthammer. Stop-transitions of Petri nets. Fundamenta Informaticae, 175(1–4):143–172, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Frei:2020:RPR

[63] Fabian Frei, Juraj Hromkovič, and Juhani Karhumäki. Roots and powers in regular languages: Recognizing nonregular properties by finite automata. Fundamenta Informaticae, 175(1–4):173–185, ???? 2020. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Genova:2020:CEM

[64] Daniela Genova, Hendrik Jan Hoogeboom, and Nataša Jonoska. Companions and an essential motion of a reaction system. Fundamenta Informaticae, 175(1–4):187–199, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Halava:2020:SWL

[65] Vesa Halava, Tero Harju, and Esa Sahla. On shuffling a word with its letter-to-letter substitution. Fundamenta Informaticae, 175(1–4):201–206, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dangalchev:2020:ACN

[66] Chavdar Dangalchev. Additional closeness and networks growth. Fundamenta Informaticae, 176(1):1–15, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jarvinen:2020:SMR

[67] Jouni Järvinen and Sándor Radeleczki. The structure of multigranular rough sets. Fundamenta Informaticae, 176(1):17–41, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2020:SUN

[68] Liucheng Liu, Wenmei Yi, Qian Yang, Hong Peng, and Jun Wang. Small universal numerical P systems with thresholds for computing functions. Fundamenta Informaticae, 176(1):43–59, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maschio:2020:TFC

[69] Samuele Maschio and Giovanni Sambin. Topology as faithful communication through relations. Fundamenta Informaticae, 176(1):61–78, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhao:2020:SNO

[70] Chenyue Zhao, Hosein Alizadeh, Behrouz Minaei, Majid Mohamadpoor, Hamid Parvin, and Mohammad Reza Mahmoudi. Social network optimization for cluster ensemble selection. Fundamenta Informaticae, 176 (1):79–102, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cheng:2020:CNI

[71] Baolei Cheng, Jianxi Fan, Qiang Lyu, Cheng-Kuan Lin, Xiaoyan Li, and Guo Chen. Constructing node-independent spanning trees in augmented cubes. Fundamenta Informaticae, 176(2):103–128, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kopczynski:2020:ARG

[72] Eryk Kopczyński. Axiomatizing rectangular grids with no extra non-unary relations. Fundamenta Informaticae, 176(2):129–138, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ghosh:2020:P

[73] Kuntal Ghosh and Sushmita Mitra. Preface. Fundamenta Informaticae, 176(2):139–140, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dasgupta:2020:PRM

[74] Abhijit Dasgupta, Losiana Nayak, Ritankar Das, Debasis Basu, Preetam Chandra, and Rajat K. De. Pattern and rule mining for identifying signatures of epileptic patients from clinical EEG data. Fundamenta Informaticae, 176(2):141–166, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Przybyszewski:2020:CDD

[75] Andrzej W. Przybyszewski, Artur Chudzik, Stanislaw Szlufik, Piotr Habela, and Dariusz M. Koziorowski. Comparison of different data mining methods to determine disease progression in dissimilar groups of Parkinson’s patients. Fundamenta Informaticae, 176(2):167–181, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chapaneri:2020:SGP

[76] Santosh Chapaneri and Deepak Jayaswal. Structured Gaussian process regression of music mood. Fundamenta Informaticae, 176(2):183–203, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Monica:2020:P

[77] Dario Della Monica, Aniello Murano, and Luigi Sauro. Preface. Fundamenta Informaticae, 176(3–4):i–ii, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Su:2020:UAN

[78] Ezgi Iraz Su. A unifying approach for nonmonotonic S4F, (reflexive) autoepistemic logic, and answer set programming. Fundamenta Informaticae, 176(3–4):205–234, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Giordano:2020:RAE

[79] Laura Giordano and Valentina Gliozzi. Reasoning about exceptions in ontologies: from the lexicographic closure to the skeptical closure. Fundamenta Informaticae, 176(3–4):235–269, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alviano:2020:UCA

[80] Mario Alviano and Carmine Dodaro. Unsatisfiable core analysis and aggregates for optimum stable model search. Fundamenta Informaticae, 176 (3–4):271–297, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tarzariol:2020:TLP

[81] Alice Tarzariol, Eugenia Zanazzo, Agostino Dovier, and Alberto Policriti. Towards a logic programming tool for cancer data analysis. Fundamenta Informaticae, 176(3–4):299–319, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alberti:2020:DOC

[82] Marco Alberti, Marco Gavanelli, Evelina Lamma, Fabrizio Riguzzi, Ken Satoh, and Riccardo Zese. Dischargeable obligations in the SCIFF framework. Fundamenta Informaticae, 176(3–4):321–348, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cantone:2020:STA

[83] Domenico Cantone, Marianna Nicolosi-Asmundo, and Daniele Francesco Santamaria. A set-theoretic approach to reasoning services for the description logic DLD4,×. Fundamenta Informaticae, 176(3–4):349–384, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Borger:2020:BTR

[84] Egon Börger and Klaus-Dieter Schewe. A behavioural theory of recursive algorithms. Fundamenta Informaticae, 177(1):1–37, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dowbor:2020:CCT

[85] Piotr Dowbor and Yan Kim. Computational classification of tubular algebras. Fundamenta Informaticae, 177(1):39–67, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ghosal:2020:PPS

[86] Purnata Ghosal and B. V. Raghavendra Rao. On proving parameterized size lower bounds for multilinear algebraic models. Fundamenta Informaticae, 177(1):69–93, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Berghammer:2020:ECL

[87] Rudolf Berghammer, Henning Schnoor, and Michael Winter. Efficient computation of the large inductive dimension using order- and graphtheoretic means. Fundamenta Informaticae, 177(2):95–113, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cho:2020:AGN

[88] Gook Hwa Cho, Seongan Lim, and Hyang-Sook Lee. Algorithms for the generalized NTRU equations and their storage analysis. Fundamenta Informaticae, 177(2):115–139, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kheirfam:2020:NPC

[89] Behrouz Kheirfam. A new predictor-corrector infeasible interior-point algorithm for linear optimization in aWide neighborhood. Fundamenta Informaticae, 177(2):141–156, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Laue:2020:SET

[90] Sören Laue, Matthias Mitterreiter, and Joachim Giesen. A simple and efficient tensor calculus for machine learning. Fundamenta Informaticae, 177(2):157–179, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:2020:DFG

[91] Cheng-Kuan Lin, Tzu-Liang Kung, Dajin Wang, and Yuan-Hsiang Teng. The diagnosability of (K_{4}–e)-free graphs under the PMC diagnosis model. Fundamenta Informaticae, 177(2):181–188, ???? 2020. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Qi:2020:CID

[92] Bin Qi, Jie Ma, and Kewei Lv. Computing interval discrete logarithm problem with restricted jump method. Fundamenta Informaticae, 177(2): 189–201, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hermenegildo:2020:P

[93] Manuel Hermenegildo, Pedro López-García, Alberto Pettorossi, and Maurizio Proietti. Preface. Fundamenta Informaticae, 177(3–4):i–iii, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Albert:2020:FRC

[94] Elvira Albert, Nikolaos Bezirgiannis, Frank de Boer, and Enrique Martin-Martin. A formal, resource consumption-preserving translation from actors with cooperative scheduling to Haskell. Fundamenta Informaticae, 177(3–4):203–234, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alpuente:2020:ACS

[95] María Alpuente, Daniel Pardo, and Alicia Villanueva. Abstract contract synthesis and verification in the symbolic K framework. Fundamenta Informaticae, 177(3–4):235–273, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bichler:2020:LRO

[96] Manuel Bichler, Michael Morak, and Stefan Woltran. lpopt: A rule optimization tool for answer set programming. Fundamenta Informaticae, 177 (3–4):275–296, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alpuente:2020:OSH

[97] María Alpuente, Angel Cuenca-Ortega, Santiago Escobar, and José Meseguer. Order-sorted homeomorphic embedding modulo combinations of associativity and/or commutativity axioms. Fundamenta Informaticae, 177(3–4):297–329, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Falaschi:2020:DSC

[98] Moreno Falaschi, Maurizio Gabbrielli, Carlos Olarte, and Catuscia Palamidessi. Dynamic slicing for concurrent constraint languages. Fundamenta Informaticae, 177(3–4):331–357, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mesnard:2020:SUC

[99] Fred Mesnard, Étienne Payet, and Germán Vidal. Selective unification in (constraint) logic programming. Fundamenta Informaticae, 177(3–4):359– 383, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tarau:2020:DES

[100] Paul Tarau. Deriving efficient sequential and parallel generators for closed simply-typed Lambda terms and normal forms. Fundamenta Informaticae, 177(3–4):385–415, ???? 2020. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hague:2021:P

[101] Matthew Hague and Igor Potapov. Preface. Fundamenta Informaticae, 178(1–2):v–vi, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bruse:2021:CMC

[102] Florian Bruse, Martin Lange, and Etienne Lozes. The complexity of model-checking tail-recursive higher-order fixpoint logic. Fundamenta Informaticae, 178(1–2):1–30, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cassez:2021:VPS

[103] Franck Cassez, Peter Gjøl Jensen, and Kim Guldstrand Larsen. Verification and parameter synthesis for real-time programs using refinement of trace abstraction. Fundamenta Informaticae, 178(1–2):31–57, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Filiot:2021:CSS

[104] Emmanuel Filiot and Pierre-Alain Reynier. Copyful streaming string transducers. Fundamenta Informaticae, 178(1–2):59–76, ???? 2021. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hutagalung:2021:TCM

[105] Milka Hutagalung. Topological characterisation of multi-buffer simulation. Fundamenta Informaticae, 178(1–2):77–99, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sproston:2021:PTA

[106] Jeremy Sproston. Probabilistic timed automata with clock-dependent probabilities. Fundamenta Informaticae, 178(1–2):101–138, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Valmari:2021:SSF

[107] Antti Valmari and Walter Vogler. Stubborn sets, frozen actions, and fair testing. Fundamenta Informaticae, 178(1–2):139–172, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Adinayev:2021:DSR

[108] Arthur Adinayev and Itamar Stein. Diamond subgraphs in the reduction graph of a one-rule string rewriting system. Fundamenta Informaticae, 178(3):173–185, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Arockiaraj:2021:OWB

[109] Micheal Arockiaraj, J. Nancy Delaila, and Jessie Abraham. Optimal wirelength of balanced complete multipartite graphs onto Cartesian product of path, cycle and trees. Fundamenta Informaticae, 178(3):187–202, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jastrzab:2021:PAM

[110] Tomasz Jastrzab, Zbigniew J. Czech, and Wojciech Wieczorek. Parallel algorithms for minimal nondeterministic finite automata inference. Fundamenta Informaticae, 178(3):203–227, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lanese:2021:CCR

[111] Ivan Lanese, Adrián Palacios, and Germán Vidal. Causal-consistent replay reversible semantics for message passing concurrent programs. Fundamenta Informaticae, 178(3):229–266, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Neethu:2021:GPN

[112] P. K. Neethu, S. V. Ullas Chandran, Manoj Changat, and Sandi Klavžar. On the general position number of complementary prisms. Fundamenta Informaticae, 178(3):267–281, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Felli:2021:P

[113] Paolo Felli, Marco Montali, and Maurizio Proietti. Preface. Fundamenta Informaticae, 178(4):i–ii, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Combi:2021:CSP

[114] Carlo Combi, Romeo Rizzi, and Pietro Sala. Checking sets of pure evolving association rules. Fundamenta Informaticae, 178(4):283–313, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cantone:2021:ISB

[115] Domenico Cantone, Marianna Nicolosi-Asmundo, and Daniele Francesco Santamaria. An improved set-based reasoner for the description logic DLD4,×. Fundamenta Informaticae, 178(4):315–346, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Conchon:2021:DPV

[116] Sylvain Conchon, Giorgio Delzanno, and Angelo Ferrando. Declarative parameterized verification of distributed protocols via the cubicle model checker. Fundamenta Informaticae, 178(4):347–378, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Benouhiba:2021:MLR

[117] Toufik Benouhiba. A multi-level refinement approach for structural synthesis of optimal probabilistic models. Fundamenta Informaticae, 179(1): 1–33, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marrakchi:2021:SSL

[118] Sirine Marrakchi and Mohamed Jemni. Static scheduling with load balancing for solving triangular band linear systems on multicore processors. Fundamenta Informaticae, 179(1):35–58, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Slapal:2021:DJC

[119] Josef Šlapal. Digital Jordan curves and surfaces with respect to a closure operator. Fundamenta Informaticae, 179(1):59–74, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Syau:2021:VPG

[120] Yu-Ru Syau, Churn-Jung Liau, and En-Bing Lin. On variable precision generalized rough sets and incomplete decision tables. Fundamenta Informaticae, 179(1):75–92, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schlingloff:2021:P

[121] H. Schlingloff and W. Penczek. Preface. Fundamenta Informaticae, 179 (2):i–ii, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2021:EEC

[122] Ludwik Czaja. Extensions of elementary cause-effect structures. Fundamenta Informaticae, 179(2):93–111, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Akili:2021:RED

[123] Samira Akili and Matthias Weidlich. Reasoning on the efficiency of distributed complex event processing. Fundamenta Informaticae, 179(2): 113–134, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yaman:2021:QVS

[124] Sinem Getir Yaman, Esteban Pavese, and Lars Grunske. Quantitative verification of stochastic regular expressions. Fundamenta Informaticae, 179(2):135–163, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grabowski:2021:ACS

[125] Adam Grabowski. Automated comparative study of some generalized rough approximations. Fundamenta Informaticae, 179(2):165–182, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Artiemjew:2021:NEM

[126] Piotr Artiemjew and Krzysztof Ropiak. A novel ensemble model — the random granular reflections. Fundamenta Informaticae, 179(2):183–203, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2021:ETB

[127] Roberto Barbuti, Pasquale Bove, Roberta Gori, Damas Gruska, Francesca Levi, and Paolo Milazzo. Encoding threshold Boolean networks into reaction systems for the analysis of gene regulatory networks. Fundamenta Informaticae, 179(2):205–225, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koprowski:2021:CSE

[128] Przemysław Koprowski. Computing singular elements modulo squares. Fundamenta Informaticae, 179(3):227–238, ???? 2021. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:2021:RAS

[129] Zhe Lin, Mihir Kumar Chakraborty, and Minghui Ma. Residuated algebraic structures in the vicinity of pre-rough algebra and decidability. Fundamenta Informaticae, 179(3):239–274, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Torres:2021:RAL

[130] Gabriela Martín Torres. On rough approximations of languages under infinite index indiscernibility relations. Fundamenta Informaticae, 179(3): 275–293, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2021:CDF

[131] Longchun Wang, Lankun Guo, and Qingguo Li. Continuous domains in formal concept analysis. Fundamenta Informaticae, 179(3):295–319, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:2021:UHL

[132] J. A. Bergstra and C. A. Middelburg. Using Hoare logic in a process algebra setting. Fundamenta Informaticae, 179(4):321–344, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Boomari:2021:RVG

[133] Hossein Boomari, Mojtaba Ostovari, and Alireza Zarei. Recognizing visibility graphs of triangulated irregular networks. Fundamenta Informaticae, 179(4):345–360, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krivka:2021:SCG

[134] Zbyněk Křivka and Alexander Meduna. Scattered context grammars with one non-context-free production are computationally complete. Fundamenta Informaticae, 179(4):361–384, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Teh:2021:FPM

[135] Wen Chean Teh, Adrian Atanasiu, and Denis C. K. Wong. Freeness problem for matrix semigroups of Parikh matrices. Fundamenta Informaticae, 179(4):385–397, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2021:NDT

[136] Zhaohao Wang. A new description of transversal matroids through rough set approach. Fundamenta Informaticae, 179(4):399–416, ???? 2021. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hirvensalo:2021:P

[137] Mika Hirvensalo, František Mráz, and Daniel Pruša. Preface. Fundamenta Informaticae, 180(1–2):v–vi, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fernau:2021:SVP

[138] Henning Fernau, Martin Kutrib, and Matthias Wendlandt. Self-verifying pushdown and queue automata. Fundamenta Informaticae, 180(1–2):1– 28, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Holzer:2021:TSS

[139] Markus Holzer, Martin Kutrib, and Friedrich Otto. Two-sided strictly locally testable languages. Fundamenta Informaticae, 180(1–2):29–51, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Malcher:2021:DQI

[140] Andreas Malcher. Decidability questions for insertion systems and related models. Fundamenta Informaticae, 180(1–2):53–76, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Otto:2021:CTR

[141] Friedrich Otto. A complete taxonomy of restarting automata without auxiliary symbols. Fundamenta Informaticae, 180(1–2):77–101, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pighizzini:2021:NSE

[142] Giovanni Pighizzini and Luca Prigioniero. Non-self-embedding grammars and descriptional complexity. Fundamenta Informaticae, 180(1–2):103– 122, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Truthe:2021:GCC

[143] Bianca Truthe. Generative capacity of contextual grammars with subregular selection languages. Fundamenta Informaticae, 180(1–2):123–150, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2021:LAW

[144] Qichao Wang. Languages accepted by weighted restarting automata. Fundamenta Informaticae, 180(1–2):151–177, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gorrieri:2021:CSB

[145] Roberto Gorrieri. Causal semantics for BPP nets with silent moves. Fundamenta Informaticae, 180(3):179–249, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Massazza:2021:DSA

[146] Paolo Massazza. A dynamical system approach to polyominoes generation. Fundamenta Informaticae, 180(3):251–273, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rudi:2021:PVA

[147] Ali Gholami Rudi. Place the vertices anywhere on the curve and simplify. Fundamenta Informaticae, 180(3):275–287, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skrzypczak:2021:P

[148] Michał Skrzypczak and Piotr Hofman. Preface. Fundamenta Informaticae, 180(4):i–ii, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Catalano:2021:LBR

[149] Costanza Catalano, Umer Azfar, Ludovic Charlier, and Raphaël M. Jungers. A linear bound on the k-rendezvous time for primitive sets of NZ matrices. Fundamenta Informaticae, 180(4):289–314, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dobronravov:2021:LSS

[150] Egor Dobronravov, Nikita Dobronravov, and Alexander Okhotin. On the length of shortest strings accepted by two-way finite automata. Fundamenta Informaticae, 180(4):315–331, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gastin:2021:RRL

[151] Paul Gastin, Amaldev Manuel, and R. Govind. Reversible regular languages: Logical and algebraic characterisations. Fundamenta Informaticae, 180(4):333–350, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kuperberg:2021:CAB

[152] Denis Kuperberg, Laureline Pinault, and Damien Pous. Coinductive algorithms for Büchi automata. Fundamenta Informaticae, 180(4):351–373, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Saarela:2021:SWL

[153] Aleksi Saarela. Separating the words of a language by counting factors. Fundamenta Informaticae, 180(4):375–393, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Casagrande:2021:P

[154] Alberto Casagrande, Eugenio G. Omodeo, and Maurizio Proietti. Preface. Fundamenta Informaticae, 181(1):v–vi, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hillston:2021:PSN

[155] Jane Hillston, Andrea Marin, Carla Piazza, and Sabina Rossi. Persistent stochastic non-interference. Fundamenta Informaticae, 181(1):1–35, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cantone:2021:CAD

[156] Domenico Cantone, Andrea De Domenico, Pietro Maugeri, and Eugenio G. Omodeo. Complexity assessments for decidable fragments of set theory. I: A taxonomy for the Boolean case. Fundamenta Informaticae, 181(1):37–69, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Costantini:2021:AMF

[157] Stefania Costantini and Andrea Formisano. Adding metalogic features to knowledge representation languages. Fundamenta Informaticae, 181(1): 71–98, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Durand-Lose:2021:P

[158] Jérôme Durand-Lose, Jarkko Kari, and Sergey Verlan. Preface. Fundamenta Informaticae, 181(2–3):i–iii, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Geffert:2021:MSC

[159] Viliam Geffert and Zuzana Bednárová. Minimal size of counters for (realtime) multicounter automata. Fundamenta Informaticae, 181(2–3):99– 127, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Whyman:2021:PCC

[160] Richard Whyman. Physical computational complexity and first-order logic. Fundamenta Informaticae, 181(2–3):129–161, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Perrot:2021:BAN

[161] Kévin Perrot, Pacôme Perrotin, and Sylvain Sené. On Boolean automata networks (de)Composition. Fundamenta Informaticae, 181(2–3):163–188, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fernau:2021:IDC

[162] Henning Fernau, Lakshmanan Kuppusamy, Rufus O. Oladele, and Indhumathi Raman. Improved descriptional complexity results for simple semiconditional grammars. Fundamenta Informaticae, 181(2–3):189–211, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nagy:2021:CIV

[163] Benedek Nagy and Sándor Vályi. Circular interval-valued computers and simulation of (red–green) Turing machines. Fundamenta Informaticae, 181 (2–3):213–238, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alhazov:2021:RBC

[164] Artiom Alhazov, Rudolf Freund, Sergiu Ivanov, and Marion Oswald. Relations between control mechanisms for sequential grammars. Fundamenta Informaticae, 181(2–3):239–271, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanderAalst:2021:FCN

[165] Wil M. P. van der Aalst. Free-choice nets with home clusters are lucent. Fundamenta Informaticae, 181(4):273–302, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Drylo:2021:CTJ

[166] Robert Dryło. Compression on the twisted Jacobi intersection. Fundamenta Informaticae, 181(4):303–312, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Perez:2021:PTC

[167] Claudia Pérez and Daniel Rivera. Polynomial-time classification of skewsymmetrizable matrices with a positive definite quasi-Cartan companion. Fundamenta Informaticae, 181(4):313–337, ???? 2021. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2021:UMD

[168] Kuize Zhang. A unified method to decentralized state detection and fault diagnosis/prediction of discrete-event systems. Fundamenta Informaticae, 181(4):339–371, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Keller:2021:P

[169] Jörg Keller and Wojciech Penczek. Preface. Fundamenta Informaticae, 182(1):v–vi, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Felli:2021:SVD

[170] Paolo Felli, Massimiliano de Leoni, and Marco Montali. Soundness verification of data-aware process models with variable-to-variable conditions. Fundamenta Informaticae, 182(1):1–29, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Andre:2021:PSA

[171] Étienne André, Emmanuel Coquard, Laurent Fribourg, Jawher Jerray, and David Lesens. Parametric schedulability analysis of a launcher flight control system under reactivity constraints. Fundamenta Informaticae, 182(1):31–67, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Andre:2021:PAA

[172] Étienne André, Didier Lime, Mathias Ramparison, and Mariëlle Stoelinga. Parametric analyses of attack-fault trees. Fundamenta Informaticae, 182 (1):69–94, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Slezak:2021:PIG

[173] Dominik Ślȩzak, Tzung-Pei Hong, and Leon S. L. Wang. Preface: Information granulation in data science and scalable computing. Fundamenta Informaticae, 182(2):i–ii, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Le:2021:KLD

[174] Linh Le, Ying Xie, and Vijay V. Raghavan. KNN loss and deep KNN. Fundamenta Informaticae, 182(2):95–110, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dagdia:2021:DSD

[175] Zaineb Chelly Dagdia and Christine Zarges. A detailed study of the distributed rough set based locality sensitive hashing feature selection technique. Fundamenta Informaticae, 182(2):111–179, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsumoto:2021:MCP

[176] Shusaku Tsumoto, Shoji Hirano, Tomohiro Kimura, and Haruko Iwata. Mining clinical process from hospital information system: a granular computing approach. Fundamenta Informaticae, 182(2):181–218, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chehreghani:2021:EAA

[177] Mostafa Haghir Chehreghani, Albert Bifet, and Talel Abdessalem. Exact and approximate algorithms for computing betweenness centrality in directed graphs. Fundamenta Informaticae, 182(3):219–242, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jin:2021:TFS

[178] Yu Jin, Bosheng Song, Yanyan Li, and Ying Zhu. Time-free solution to independent set problem using P systems with active membranes. Fundamenta Informaticae, 182(3):243–255, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2021:EAM

[179] Viet Dung Nguyen, Ba Thai Pham, and Phan Thuan Do. Efficient algorithms for maximum induced matching problem in permutation and trapezoid graphs. Fundamenta Informaticae, 182(3):257–283, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sujana:2021:EFB

[180] G. Jessy Sujana, T. M. Rajalaxmi, Indra Rajasingh, and R. Sundara Rajan. Edge forcing in butterfly networks. Fundamenta Informaticae, 182(3):285–299, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zarrabi:2021:QPV

[181] Mohammad Reza Zarrabi and Nasrollah Moghaddam Charkari. Querypoints visibility constraint minimum link paths in simple polygons. Fundamenta Informaticae, 182(3):301–319, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhengwei:2021:TBF

[182] Xie Zhengwei, Qiu Daowen, Cai Guangya, Jozef Gruska, and Paulo Mateus. Testing Boolean functions properties. Fundamenta Informaticae, 182(4):321–344, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sureson:2021:IAF

[183] Claude Sureson. The inverse of Ackermann function is computable in linear time. Fundamenta Informaticae, 182(4):345–361, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wehler:2021:PFC

[184] Joachim Wehler. Perpetual free-choice Petri nets are lucent proof of a theorem of van der aalst using CP -exhaustions. Fundamenta Informaticae, 182(4):363–393, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Donatelli:2021:P

[185] Susanna Donatelli, Stefan Haar, and Slawomir Lasota. Preface. Fundamenta Informaticae, 183(1–2):v–vi, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Devillers:2021:APT

[186] Raymond Devillers. Articulations and products of transition systems and their applications to Petri net synthesis. Fundamenta Informaticae, 183 (1–2):1–31, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Finkel:2021:CTF

[187] Alain Finkel, Serge Haddad, and Igor Khmelnitsky. Coverability, termination, and finiteness in recursive Petri nets. Fundamenta Informaticae, 183(1–2):33–66, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Escrig:2021:IRS

[188] David de Frutos Escrig, Maciej Koutny, and Łukasz Mikulski. Investigating reversibility of steps in Petri nets. Fundamenta Informaticae, 183 (1–2):67–96, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lime:2021:CPP

[189] Didier Lime, Olivier H. Roux, and Charlotte Seidner. Cost problems for parametric time Petri nets. Fundamenta Informaticae, 183(1–2):97– 123, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tredup:2021:CSB

[190] Ronny Tredup. The complexity of synthesis of b-bounded Petri nets. Fundamenta Informaticae, 183(1–2):125–167, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:2021:P

[191] Ryszard Janicki, Slawomir Lasota, and Natalia Sidorova. Preface. Fundamenta Informaticae, 183(3–4):i–ii, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Allamigeon:2021:PAD

[192] Xavier Allamigeon, Marin Boyet, and Stéphane Gaubert. Piecewise affine dynamical models of Petri nets — application to emergency call centers. Fundamenta Informaticae, 183(3–4):169–201, ???? 2021. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fahland:2021:IUE

[193] Dirk Fahland, Vadim Denisov, and Wil. M. P. van der Aalst. Inferring unobserved events in systems with shared resources and queues. Fundamenta Informaticae, 183(3–4):203–242, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Finkel:2021:EPN

[194] Olivier Finkel and Michał Skrzypczak. On the expressive power of nondeterministic and unambiguous Petri nets over infinite words. Fundamenta Informaticae, 183(3–4):243–291, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kalenkova:2021:ARP

[195] Anna Kalenkova, Josep Carmona, Artem Polyvyanyy, and Marcello La Rosa. Automated repair of process models with non-local constraints using state-based region theory. Fundamenta Informaticae, 183(3–4):293– 317, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Thierry-Mieg:2021:SSM

[196] Yann Thierry-Mieg. Symbolic and structural model-checking. Fundamenta Informaticae, 183(3–4):319–342, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ganty:2021:CBP

[197] Pierre Ganty, Elena Gutiérrez, and Pedro Valero. A congruence-based perspective on finite tree automata. Fundamenta Informaticae, 184(1): 1–47, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gonzalez:2021:GTF

[198] Jesús Arturo Jiménez González. A graph theoretical framework for the strong Gram classification of non-negative unit forms of Dynkin type A_{n}. Fundamenta Informaticae, 184(1):49–82, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fajardo:2021:RBA

[199] William Fajardo. Right Buchberger algorithm over bijective skew PBW extensions. Fundamenta Informaticae, 184(2):83–105, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wronski:2021:HDC

[200] Michał Wroński, Tomasz Kijko, and Robert Dryło. High-degree compression functions on alternative models of elliptic curves and their applications. Fundamenta Informaticae, 184(2):107–139, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zielinski:2021:NDM

[201] Bartosz Zieliński. A non-deterministic multiset query language. Fundamenta Informaticae, 184(2):141–180, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Adamson:2021:HEM

[202] Duncan Adamson, Argyrios Deligkas, Vladimir Gusev, and Igor Potapov. On the hardness of energy minimisation for crystal structure prediction. Fundamenta Informaticae, 184(3):181–203, ???? 2021. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Keikha:2021:CGM

[203] Vahideh Keikha, Sepideh Aghamolaei, Ali Mohades, and Mohammad Ghodsi. Clustering geometrically-modeled points in the aggregated uncertainty model. Fundamenta Informaticae, 184(3):205–231, ???? 2021. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mitrovic:2021:TCS

[204] Melanija Mitrović, Mahouton Norbert Hounkonnou, and Marian Alexandru Baroni. Theory of constructive semigroups with apartness — foundations, development and practice. Fundamenta Informaticae, 184(3):233– 271, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barylska:2021:ACR

[205] Kamila Barylska and Anna Gogolińska. Acyclic and cyclic reversing computations in Petri nets. Fundamenta Informaticae, 184(4):273–296, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Darkey-Mensah:2021:CLS

[206] Mawunyo Kofi Darkey-Mensah and Beata Rothkegel. Computing the length of sum of squares and Pythagoras element in a global field. Fundamenta Informaticae, 184(4):297–306, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sima:2021:PTC

[207] Jiří Šíma and Stanislav Žák. A polynomial-time construction of a hitting set for read-once branching programs of width 3. Fundamenta Informaticae, 184(4):307–354, ???? 2021. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dolce:2022:MPP

[208] Francesco Dolce and Edita Pelantová. On morphisms preserving palindromic richness. Fundamenta Informaticae, 185(1):1–25, ???? 2022. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fates:2022:SSC

[209] Nazim Fatès, Irène Marcovici, and Siamak Taati. Self-stabilisation of cellular automata on tilings. Fundamenta Informaticae, 185(1):27–82, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Forcan:2022:SSW

[210] Jovana Forcan and Mirjana Mikalački. Spanning structures in walker– breaker games. Fundamenta Informaticae, 185(1):83–97, ???? 2022. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baca:2022:LAV

[211] Martin Bača, Andrea Semaničová-Feňovčíková, Ruei-Ting Lai, and Tao-Ming Wang. On local antimagic vertex coloring for complete full t-ary trees. Fundamenta Informaticae, 185(2):99–113, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Danvy:2022:GTB

[212] Olivier Danvy. Getting there and back again. Fundamenta Informaticae, 185(2):115–183, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Garzon:2022:DNC

[213] Ester M. Garzón, José A. Martínez, Juan J. Moreno, and María L. Puertas. On the 2-domination number of cylinders with small cycles. Fundamenta Informaticae, 185(2):185–199, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martinez:2022:PDR

[214] A. Cabrera Martínez, C. García-Gómez, and J. A. Rodríguez-Velázquez. Perfect domination, Roman domination and perfect Roman domination in lexicographic product graphs. Fundamenta Informaticae, 185(3):201– 220, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gonzalez:2022:CIN

[215] Jesús Arturo Jiménez González. Coxeter invariants for non-negative unit forms of Dynkin type A_{r}. Fundamenta Informaticae, 185(3):221–246, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schmidt-Schauss:2022:NUM

[216] Manfred Schmidt-Schauß, Temur Kutsia, Jordi Levy, Mateu Villaret, and Yunus Kutz. Nominal unification and matching of higher order expressions with recursive let. Fundamenta Informaticae, 185(3):247–283, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abhishek:2022:RNG

[217] Kunal Abhishek and E. George Dharma Prakash Raj. On random number generation for kernel applications. Fundamenta Informaticae, 185(4):285– 311, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Geuvers:2022:CBE

[218] Herman Geuvers and Rob Nederpelt. Characteristics of de Bruijn’s early proof checker Automath. Fundamenta Informaticae, 185(4):313–336, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kutrib:2022:CDP

[219] Martin Kutrib, Andreas Malcher, Carlo Mereghetti, and Beatrice Palano. Computational and descriptional power of nondeterministic iterated uniform finite-state transducers. Fundamenta Informaticae, 185(4):337–356, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:2022:LDA

[220] Cheng-Kuan Lin, Tzu-Liang Kung, Chun-Nan Hung, and Yuan-Hsiang Teng. A local diagnosis algorithm for hypercube-like networks under the BGM diagnosis model. Fundamenta Informaticae, 185(4):357–373, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Avron:2022:P

[221] Arnon Avron, Nachum Dershowitz, and Alexander Rabinovich. Preface. Fundamenta Informaticae, 186(1–4):v–viii, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abramsky:2022:SPE

[222] Samson Abramsky. Structure and power: an emerging landscape. Fundamenta Informaticae, 186(1–4):1–26, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Arnold:2022:ACS

[223] André Arnold, Patrick Cégielski, and Irène Guessarian. Affine completeness of some free binary algebras. Fundamenta Informaticae, 186(1–4): 27–44, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Artemov:2022:TSE

[224] Sergei Artemov. Towards syntactic epistemic logic. Fundamenta Informaticae, 186(1–4):45–62, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brutsch:2022:SIG

[225] Benedikt Brütsch and Wolfgang Thomas. Solving infinite games in the Baire space. Fundamenta Informaticae, 186(1–4):63–88, ???? 2022. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Courcelle:2022:OTT

[226] Bruno Courcelle. Order-theoretic trees: Monadic second-order descriptions and regularity. Fundamenta Informaticae, 186(1–4):89–120, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Francez:2022:GFF

[227] Nissim Francez. A generalization of falsity in finitely-many valued logics. Fundamenta Informaticae, 186(1–4):121–132, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gurevich:2022:ICM

[228] Yuri Gurevich. The 1966 International Congress of Mathematicians: a micro-memoir. Fundamenta Informaticae, 186(1–4):133–141, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kaminski:2022:NCN

[229] Michael Kaminski. A note on calculi for non-deterministic many-valued logics. Fundamenta Informaticae, 186(1–4):143–153, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kotek:2022:TMP

[230] Tomer Kotek and Johann A. Makowsky. On the Tutte and matching polynomials for complete graphs. Fundamenta Informaticae, 186(1–4): 155–173, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kasjan:2022:DSO

[231] Stanisław Kasjan and Damian Niwiński. Daniel Simson obituary. Fundamenta Informaticae, 187(1):v–vii, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bhagat:2022:GMN

[232] Subhash Bhagat, Abhinav Chakraborty, Bibhuti Das, and Krishnendu Mukhopadhyaya. Gathering over meeting nodes in infinite grid. Fundamenta Informaticae, 187(1):1–30, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redeker:2022:NCP

[233] Markus Redeker. Number conservation via particle flow in onedimensional cellular automata. Fundamenta Informaticae, 187(1):31–59, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhou:2022:NGF

[234] Sizhong Zhou. A note of generalization of fractional ID-factor-critical graphs. Fundamenta Informaticae, 187(1):61–69, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Buchs:2022:P

[235] Didier Buchs, Josep Carmona, and Jetty Kleijn. Preface. Fundamenta Informaticae, 187(2–4):i–iii, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abbes:2022:IPC

[236] Samy Abbes. Introduction to probabilistic concurrent systems. Fundamenta Informaticae, 187(2–4):71–102, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amat:2022:PAP

[237] Nicolas Amat, Bernard Berthomieu, and Silvano Dal Zilio. A polyhedral abstraction for Petri nets and its application to SMT-based model checking. Fundamenta Informaticae, 187(2–4):103–138, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Devillers:2022:SPI

[238] Raymond Devillers and Ronny Tredup. Synthesis of pure and impure Petri nets with restricted place-environments: Complexity issues. Fundamenta Informaticae, 187(2–4):139–165, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Devillers:2022:SBT

[239] Raymond Devillers and Ronny Tredup. Some basic techniques allowing Petri net synthesis: Complexity and algorithmic issues. Fundamenta Informaticae, 187(2–4):167–196, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esparza:2022:CPI

[240] Javier Esparza, Mikhail Raskin, and Christoph Welzel. Computing parameterized invariants of parameterized Petri nets. Fundamenta Informaticae, 187(2–4):197–243, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wallner:2022:SAU

[241] Sophie Wallner and Karsten Wolf. Skeleton abstraction for universal temporal properties. Fundamenta Informaticae, 187(2–4):245–272, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gh:2022:FHC

[242] Behrooz Bagheri Gh., Tomas Feder, Herbert Fleischner, and Carlos Subi. On finding Hamiltonian cycles in barnette graphs. Fundamenta Informaticae, 188(1):1–14, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bes:2022:DDI

[243] Alexis Bès and Christian Choffrut. Decidability of definability issues in the theory of real addition. Fundamenta Informaticae, 188(1):15–39, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mrozek:2022:UAS

[244] Ireneusz Mrozek, Nikolai A. Shevchenko, and Vyacheslav N. Yarmolik. Universal address sequence generator for memory built-in self-test. Fundamenta Informaticae, 188(1):41–61, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eberbach:2022:CCM

[245] Eugene Eberbach. On completeness of cost metrics and meta-search algorithms in $-calculus. Fundamenta Informaticae, 188(2):63–90, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lee:2022:IUB

[246] Hyang-Sook Lee, Seongan Lim, Ikkwon Yie, and Aaram Yun. On insecure uses of BGN for privacy preserving data aggregation protocols. Fundamenta Informaticae, 188(2):91–101, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Macyna:2022:AMP

[247] Wojciech Macyna and Michal Kukowski. Adaptive merging on phase change memory. Fundamenta Informaticae, 188(2):103–126, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barozzini:2022:CAS

[248] David Barozzini, Lorenzo Clemente, Thomas Colcombet, and Paweł Parys. Cost automata, safe schemes, and downward closures. Fundamenta Informaticae, 188(3):127–178, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jancar:2022:SLI

[249] Petr Jančar and Jiří Valušek. Structural liveness of immediate observation Petri nets. Fundamenta Informaticae, 188(3):179–215, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).